# On the Convergence of Jacobian-Free Backpropagation for Optimal Control Problems with Implicit Hamiltonians

### v1

@@ Appendix B Problem Formulation for Optimal Consumption Savings
We consider a model for individual’s consumption optimization. The wealth $x_{t}$, $t\in[0,T]$ of an individual follows
@@
where $u_{i}(t)>h_{i}(t)$ for $i=1,\ldots,m$, and $\gamma\neq 1$, $\epsilon$, $\delta$, $\eta$, $\theta$ are positive constants. The consumption behavior exhibits habit formation, that is, the current utility depends on the consumption relative to an individual’s habit level established from past consumption. Then, the Hamiltonian is
@@
Let $\frac{\partial\mathcal{H}}{\partial u}=0$, then $e^{-\delta t}(u-h)^{-\gamma}-p_{x}+\alpha\eta p_{h}u^{\eta-1}=0$. This first-order condition does not have a closed-form solution in general.
@@
It is possible to prove convergence in expectation of JFB-based SGD using the iteration  (22) to a neighborhood of a critial point of $J$ under weaker assumptions than Theorem  4.10 depends on. This section was inspired by (Demidovich et al., 2024).
@@
in the proof of Lemma  4.7 still holds. Applying Assumption  C.1 to this equation gives
@@
Under these assumptions, everything in the proof of Lemma  4.8 up to and including the line
@@
still holds. Replacing $F$ with $\mathbb{E}_{x}[J_{x}]$, $\mathbb{E}_{\xi_{j}}[d_{\xi_{j}}^{JFB}(\theta_{j})]$ with $\mathbb{E}_{x}[d_{x}^{JFB}(\theta_{j})]$ like in Lemma  4.9, and applying the result of Lemma  C.3 and the upper bound $\mathbb{E}_{\xi_{j}}[\|d_{\xi_{j}}^{JFB}(\theta)\|^{2}]\leq\frac{B_{max}^{2}T^{2}}{\beta}$, we have
@@
$\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{2})]]-\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{1})]]$ $\leq-\alpha(\epsilon_{v}\mathbb{E}[\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{1})]\|^{2}]-\epsilon_{1}^{\prime})+\frac{\alpha^{2}}{2}L_{J}\left(\frac{B_{max}^{2}T^{2}}{\beta}\right)$
$\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{2})]]$ $\leq\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{1})]]-\alpha(\epsilon_{v}\mathbb{E}[\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{1})]\|^{2}]-\epsilon_{1}^{\prime})+\frac{\alpha^{2}}{2}L_{J}\left(\frac{B_{max}^{2}T^{2}}{\beta}\right)$
$\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{2})]]$ $\leq\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{0})]]-\alpha(\epsilon_{v}\mathbb{E}[\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{0})]\|^{2}]-\epsilon_{1}^{\prime})+\frac{\alpha^{2}}{2}L_{J}\left(\frac{B_{max}^{2}T^{2}}{\beta}\right)$-
$\alpha(\epsilon_{v}\mathbb{E}[\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{1})]\|^{2}]-\epsilon_{1}^{\prime})+\frac{\alpha^{2}}{2}L_{J}\left(\frac{B_{max}^{2}T^{2}}{\beta}\right)$
$\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{2})]]-\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{0})]]$ $\leq-\alpha\sum_{j=0}^{1}(\epsilon_{v}\mathbb{E}[\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\|^{2}]-\epsilon_{1}^{\prime})+(2)\frac{\alpha^{2}}{2}L_{J}\left(\frac{B_{max}^{2}T^{2}}{\beta}\right)$
$\vdots$
$\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{K})]]-\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{0})]]$ $\leq-\alpha\sum_{j=0}^{K-1}(\epsilon_{v}\mathbb{E}[\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\|^{2}]-\epsilon_{1}^{\prime})+K\left(\frac{\alpha^{2}L_{J}B_{max}^{2}T}{2\beta}\right)$
Because $J_{x}(\theta)$ is bounded from below by $J_{inf}$ (restrict domain if necessary), the last line above becomes, using linearity of expectation,

### v2

@@ Appendix B Problem Formulation for Optimal Consumption Savings
We consider the consumption optimization of multiple investors. This is adapted from (Detemple & Zapatero, 1992; Angoshtari et al., 2023). The wealth $x_{t}$, $t\in[0,T]$ of an individual follows
@@
where $u_{i}(t)>h_{i}(t)$ for $i=1,\ldots,m$, and $\gamma\neq 1$, $\epsilon$, $\delta$, $\eta$, $\theta$ are positive constants. We use a reparametrization to ensure the constraint $u_{i}(t)-h_{i}(t)>0$ for all $i$ in experiments. The consumption behavior exhibits habit formation, that is, the current utility depends on the consumption relative to an individual’s habit level established from past consumption. Then, the Hamiltonian is
@@
Let $\frac{\partial\mathcal{H}}{\partial u}=0$, then $e^{-\delta t}(u-h)^{-\gamma}-p_{x}+\alpha\eta p_{h}u^{\eta-1}=0$. Except when $\eta=\theta=1$, this first-order condition does not have a closed-form solution in general.
@@
It is possible to prove convergence in expectation of JFB-based SGD using the iteration  (22) to a neighborhood of a critical point of $J$ under weaker assumptions than Theorem  4.11 depends on. This section was inspired by (Demidovich et al., 2024).
@@
in the proof of Lemma  4.8 still holds. Applying Assumption  C.1 to this equation gives
@@
Under these assumptions, everything in the proof of Lemma  4.10 up to and including the line
@@
still holds. Replacing $F$ with $\mathbb{E}_{x}[J_{x}]$, $\mathbb{E}_{\xi_{j}}[d_{\xi_{j}}^{JFB}(\theta_{j})]$ with $\mathbb{E}_{x}[d_{x}^{JFB}(\theta_{j})]$ like in Lemma  4.10, and applying the result of Lemma  C.3 and the upper bound $\mathbb{E}_{\xi_{j}}[\|d_{\xi_{j}}^{JFB}(\theta)\|^{2}]\leq\frac{B_{max}^{2}T^{2}}{\beta}$, we have
@@
$\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{2})]]-\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{1})]]$ $\leq-\alpha(\epsilon_{v}\mathbb{E}[\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{1})]\|^{2}]-\epsilon_{1}^{\prime})+\frac{\alpha^{2}}{2}L_{J}\left(\frac{B_{max}^{2}T^{2}}{\beta}\right)$ (85)
$\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{2})]]$ $\leq\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{1})]]-\alpha(\epsilon_{v}\mathbb{E}[\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{1})]\|^{2}]-\epsilon_{1}^{\prime})+\frac{\alpha^{2}}{2}L_{J}\left(\frac{B_{max}^{2}T^{2}}{\beta}\right)$ (86)
$\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{2})]]$ $\leq\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{0})]]-\alpha(\epsilon_{v}\mathbb{E}[\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{0})]\|^{2}]-\epsilon_{1}^{\prime})+\frac{\alpha^{2}}{2}L_{J}\left(\frac{B_{max}^{2}T^{2}}{\beta}\right)$- (87)
$\alpha(\epsilon_{v}\mathbb{E}[\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{1})]\|^{2}]-\epsilon_{1}^{\prime})+\frac{\alpha^{2}}{2}L_{J}\left(\frac{B_{max}^{2}T^{2}}{\beta}\right)$ (88)
$\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{2})]]-\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{0})]]$ $\leq-\alpha\sum_{j=0}^{1}(\epsilon_{v}\mathbb{E}[\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\|^{2}]-\epsilon_{1}^{\prime})+(2)\frac{\alpha^{2}}{2}L_{J}\left(\frac{B_{max}^{2}T^{2}}{\beta}\right)$ (89)
$\vdots$ (90)
$\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{K})]]-\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{0})]]$ $\leq-\alpha\sum_{j=0}^{K-1}(\epsilon_{v}\mathbb{E}[\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\|^{2}]-\epsilon_{1}^{\prime})+K\left(\frac{\alpha^{2}L_{J}B_{max}^{2}T}{2\beta}\right)$ (91)
Because $J_{x}(\theta)$ is bounded from below by $J_{inf}$ (restrict domain if necessary), (91) becomes, using linearity of expectation,
@@
Appendix D Convergence to Critical Point Under Slightly Stronger Version of Assumption  C.1
With a slight modification to Assumption  C.1, it is possible to prove the convergence of the iteration  (22) to a critical point, not just to a neighborhood. This assumption is marginally stronger than Assumption  C.1 but is significantly weaker than Assumption  4.5.
Assumption D.1.
Modify Assumption  4.5 to be $\exists 0<\delta_{var}<\lambda_{-}-\gamma\lambda_{+}$ and a sequence of $\{\theta_{j}\}_{j=0}^{\infty}$ such that, for a nonnegative sequence $\{\epsilon_{j}\}_{j=0}^{\infty}$ that satisfies $\sum_{j=0}^{\infty}\epsilon_{j}<\infty$, and $\forall z,u,j$
$\text{max}(\text{Var}_{x}[v_{\theta_{j},x}(t)],\text{Var}_{x}[w_{\theta_{j},x}(t)])^{2}\leq\epsilon_{j}+\delta_{var}\|\mathbb{E}_{x}[M_{\theta_{j}}v_{\theta_{j},x}]\|^{2}$
Lemma D.2.
Under the Assumptions of  4.14.4 and Assumption  D.1, $\forall z,u,j$
$\langle\mathbb{E}_{x}[v_{\theta_{j},x}(t)],\mathbb{E}_{x}[w_{\theta_{j},x}(t)]\rangle\geq\delta_{\theta_{j}}^{2}-\epsilon_{j}$ (95)
where $\delta_{\theta}$ is defined in Assumption  4.6 and $\epsilon_{j}$ is defined in Assumption  D.1
The proof of this lemma is essentially identical to that of Lemma  C.2 and is thus omitted.
To prove convergence under Assumption  D.1, we will need the following lemma, whose proof is omitted because it is essentially identical to that of Lemma  C.3.
Lemma D.3.
Under Assumptions  4.14.4,  D.1, and  4.6, $\forall u,z,j$
$\mathbb{E}_{x}[\nabla_{\theta_{j}}J_{x}]^{\top}\mathbb{E}_{x}[d_{x}^{JFB}]\geq\epsilon_{v}\|\mathbb{E}_{x}[\nabla_{\theta_{j}}J_{x}]\|^{2}-\epsilon_{j}^{\prime}$
where $\epsilon_{v}$ is given in Assumption  4.6 and $\epsilon_{j}^{\prime}=T^{2}\epsilon_{j}$ with $\epsilon_{j}$ is from Assumption  D.1
We can now prove convergence to the critical point under Assumption  D.1.
Theorem D.4.
Suppose the hypotheses of Lemma  D.3 are true with the sequence $\{\epsilon_{j}\}_{j=0}^{\infty}$ from Assumption  D.1 satisfying $\epsilon_{0}<\frac{\epsilon_{v}B_{max}^{2}}{\beta(1-\gamma)^{2}}$. Then, the JFB-based SGD iteration  (22) with decreasing step size/learning rate $\alpha_{j}>0,\sum_{j=0}^{\infty}\alpha_{j}=\infty,\sum_{j=0}^{\infty}\alpha_{j}^{2}<\infty$, and $\alpha_{j}\leq\frac{2\epsilon_{v}}{L_{J}(1-\gamma)^{2}}-\frac{2\epsilon_{j}^{\prime}\beta}{L_{J}B_{max}^{2}T^{2}}\forall j$ will converge in (total) expectation to a critical point. In other words,
$\lim_{K\rightarrow\infty}\mathbb{E}\left[\frac{1}{K}\sum_{j=1}^{K}\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}]\|^{2}\right]=0$ (96)
Proof.
To prove this theorem, an approach identical to the beginning of the proof of Theorem  C.4 yields the expression below $\forall j$
$\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{j+1})]]-\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{j})]]$ $\leq-\alpha_{j}(\epsilon_{v}\mathbb{E}[\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\|^{2}]-\epsilon_{j}^{\prime})+\frac{\alpha_{j}^{2}}{2}L_{J}\left(\frac{B_{max}^{2}T^{2}}{\beta}\right)$ (97)
$\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{j+1})]]-\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{j})]]$ $\leq-\alpha_{j}\epsilon_{v}\mathbb{E}[\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\|^{2}]+\alpha_{j}\epsilon_{j}^{\prime}+\frac{\alpha_{j}^{2}}{2}L_{J}\left(\frac{B_{max}^{2}T^{2}}{\beta}\right)$ (98)
For the RHS of (98) to be $\leq 0$, it is necessary that
$-\alpha_{j}\epsilon_{v}\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\|^{2}+\alpha_{j}\epsilon_{j}^{\prime}+\frac{\alpha_{j}^{2}L_{J}B_{max}^{2}T^{2}}{2\beta}$ $\leq 0$ (99)
$\frac{\alpha_{j}^{2}L_{J}B_{max}^{2}T^{2}}{2\beta}$ $\leq\alpha_{j}\epsilon_{v}\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\|^{2}-\alpha_{j}\epsilon_{j}^{\prime}$ (100)
$\frac{\alpha_{j}L_{J}B_{max}^{2}T^{2}}{2\beta}$ $\leq\epsilon_{v}\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\|^{2}-\epsilon_{j}^{\prime}$ (101)
$\frac{\alpha_{j}L_{J}B_{max}^{2}T^{2}}{2\beta}$ $\leq\epsilon_{v}\left(\frac{B_{max}^{2}T^{2}}{\beta(1-\gamma)^{2}}\right)-\epsilon_{j}^{\prime}$ (102)
$\alpha_{j}$ $\leq\frac{2}{L_{J}}\left(\frac{\epsilon_{v}}{(1-\gamma)^{2}}-\frac{\epsilon_{j}^{\prime}\beta}{B_{max}^{2}T^{2}}\right)$ (103)
as was assumed. For the RHS of (103) to be >0>0, it is necessary that
$0$ $<\frac{2\epsilon_{v}}{L_{J}(1-\gamma)^{2}}-\frac{2\epsilon_{j}^{\prime}\beta}{L_{j}B_{max}^{2}T^{2}}$ (104)
$\epsilon_{j}^{\prime}$ $<\frac{\epsilon_{v}B_{max}^{2}T^{2}}{\beta(1-\gamma)^{2}}$ (105)
$\epsilon_{j}$ $<\frac{\epsilon_{v}B_{max}^{2}}{\beta(1-\gamma)^{2}}$ (106)
Since $\epsilon_{j}^{\prime}=T^{2}\epsilon_{j}$. Thus, descent toward the critical point has been proved for each step $j$. Then, using an argument identical to that of (86)-(91), it follows that, $\forall j$,
$\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{K})]]-\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{0})]]\leq-\epsilon_{v}\sum_{j=0}^{K-1}\alpha_{j}\mathbb{E}[\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\|^{2}]+\sum_{j=0}^{K-1}\alpha_{j}\epsilon_{j}^{\prime}+\left(\frac{L_{J}B_{max}^{2}T}{2\beta}\right)\sum_{j=0}^{K-1}\alpha_{j}^{2}$ (107)
Because $J_{x}$ is bounded from below by $J_{inf}$ by Assumption  4.1, algebraically rearranging the last line above yields, with $A_{K}=\sum_{j=0}^{K-1}\alpha_{j}$,
$\sum_{j=0}^{K-1}\alpha_{j}\mathbb{E}[\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\|^{2}]\leq\frac{\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{0})]]-J_{inf}}{\epsilon_{v}}+\frac{1}{\epsilon_{v}}\sum_{j=0}^{K-1}\alpha_{j}\epsilon_{j}^{\prime}+\frac{L_{J}B_{max}^{2}T^{2}}{2\beta\epsilon_{v}}\sum_{j=0}^{K-1}\alpha_{j}^{2}$ (108)
$\frac{1}{A_{K}}\sum_{j=0}^{K-1}\alpha_{j}\mathbb{E}[\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\|^{2}]\leq\frac{\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{0})]]-J_{inf}}{\epsilon_{v}A_{K}}+\frac{1}{\epsilon_{v}A_{K}}\sum_{j=0}^{K-1}\alpha_{j}\epsilon_{j}^{\prime}+\frac{L_{J}B_{max}^{2}T}{2\beta\epsilon_{v}A_{K}}\sum_{j=0}^{K-1}\alpha_{j}^{2}$ (109)
Because $\sum_{j=0}^{\infty}\epsilon_{j}<\infty$ it follows that the partial sums of $\{\epsilon_{j}^{\prime}\}_{j=0}^{\infty}$ are bounded. Because $\{\alpha_{j}\}_{j=0}^{\infty}$ is a decreasing sequence, it follows by the Dirichlet Test that the product series $\sum_{j=0}^{\infty}\alpha_{j}\epsilon_{1}^{\prime}$ converges. Thus, using linearity of expectation,  (109) becomes
$\mathbb{E}\left[\frac{1}{A_{K}}\sum_{j=0}^{K-1}\alpha_{j}\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\|^{2}\right]\leq\frac{\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{0})]]-J_{inf}}{\epsilon_{v}A_{K}}+\frac{1}{\epsilon_{v}A_{K}}\sum_{j=0}^{K-1}\alpha_{j}\epsilon_{j}^{\prime}+\frac{L_{J}B_{max}^{2}T^{2}}{2\beta\epsilon_{v}A_{K}}\sum_{j=0}^{K-1}\alpha_{j}^{2}$ (110)
Hence, since $\lim_{K\rightarrow\infty}A_{K}=\infty$, we have
$\leq\lim_{K\rightarrow\infty}\left[\frac{2\beta(\mathbb{E}[\mathbb{E}_{x}[J_{x}(\theta_{0})]]-J_{inf})+2\beta\sum_{j=0}^{K-1}\alpha_{j}\epsilon_{j}^{\prime}+L_{J}B_{max}^{2}T^{2}{\sum_{j=0}^{K-1}\alpha_{j}^{2}}}{2\beta\epsilon_{v}A_{K}}\right]$
$=0$
Under the assumptions of Theorem  D.4, it is possible to prove Theorem  4.12 and Corollary  4.13. These proofs are omitted because they are identical to those already shown.